OpenStudy (anonymous):

Write y = –2/3(Two thirds) + 7 in standard form using integers. A. 2x + 3y = 21 B. 3x – 2y = 21 C.–2x – 3y = 21 D.–2x + 3y = 7 @LogicalApple I Promise I'll stop bugging you soon.

OpenStudy (anonymous):

y = -2/3 x + 7 To write in standard form, first get rid of all the fractions. In this case there is only one fraction to get rid of. To get rid of it, multiply everything by 3. Once you do that, then bring the variables over on one side of the equation.

OpenStudy (anonymous):

–2x – 3y = 21, is this correct? @LogicalApple

OpenStudy (anonymous):

Almost. If we multiply everything by 3 we obtain: 3y = -2x + 21 What happens when we add 2x to both sides?

OpenStudy (anonymous):

I don't know, sadly, could you explain?

OpenStudy (anonymous):

If we add 2x to both sides then we get: 3y + 2x = -2x + 2x + 21 3y + 2x = 21 And this is the same as A

OpenStudy (anonymous):

I don't really understand adding 2 to both sides got that. Or how you even added 2 to both sides.

OpenStudy (anonymous):

We didn't add 2, we added "2x". I did this so that the "-2x" will cancel on the right side, leaving 0 + 21. But what we do to one side we do to the other. So we add +2x to the left side. This gives us 3y + 2x on the left side and 21 on the right side.

OpenStudy (anonymous):

@LogicalApple Here is another question exactly the same as this, but without fractions. Write y = –0.5x + 0.8 in standard form using integers. Would the answer be 5x + 10y = 8?

OpenStudy (anonymous):

y = -0.5x + 0.8 Multiply everything by 10 10y = -5x + 8 Move the variables to the other side 5x + 10y = 8 And you get the same answer :) You got it.